Computers manufactured by a certain company have a serial number consisting of a letter of the alphabet followed by a four-digit number. If all the serial numbers of this type have been used, how many sets have already been manufactured?
Answer by jim_thompson5910(35256)

26*10^4 = 26*10000 = 26,000 computers have already been manufactured.